A boy of mass 60kg jumps of a building of height 3.2m. Calculate the time and velocity with which he would fall to the ground. Also calculate the acceleration with which he would fall if his weight was reduced to how much he would have weighed if he was on the moon. (g = 10m/s2)


1
Expert's answer
2022-01-31T10:13:54-0500

The time and velocity:


"t=\sqrt\dfrac{2h}g=0.81\text{ s},\\\space\\
v=\sqrt{2gh}=7.9\text{ m/s}."


No matter how much the boy weights, on planet Earth, he would fall at 9.8 m/s2.